OpenStudy (anonymous):

-4x+5(x-2)=-16+10 What is the solution to the equation?

OpenStudy (anonymous):

What do you think the answer is? I've assisted you with two, try this one yourself.

OpenStudy (anonymous):

um how about idk i 've tried and i dont get the right answer

OpenStudy (anonymous):

Shoot, I gtg. Maybe @Vocaloid can help you?

OpenStudy (anonymous):

thanks

Vocaloid (vocaloid):

start by distributing 5(x-2) = ?

OpenStudy (anonymous):

(5x-10)

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

good, now lets keep going -4x + 5x - 10 = -16 + 10 now add 10 to boh sides

OpenStudy (anonymous):

ok im not good at this i know -16 +10 will = -6 but on the other side i dont get

Vocaloid (vocaloid):

-4x + 5x = -16 + 10 + 10 x = 4

OpenStudy (anonymous):

oooooh